# Multi-Agent DRL for Multi-Objective Twin Migration Routing with Workload Prediction in 6G-enabled IoV

### v1

@@ VI-B Algorithm Design
Algorithm 2 illustrates the pseudocode of the DM-MAPPO algorithm, whose computational complexity is $\mathcal{O}\left(End\times V\times t_{v}^{end}\times(M+end)\right)$, where $End$ represents the number of complete interactions of all agents with the environment, and $end$ represents the training epochs of each actor network.
@@ VII-A Parameter Settings
We first present the parameter settings for the experiments in the VT migration routing scenario. The scenario is sampled every $30\>\rm{s}$. All UAVs patrol among RSUs at a speed of $3.7\>\rm{m/s}$. For the network model, the mean of additive Gaussian white noise $\sigma$ is set to 00, with a variance of $10$ [8]. The vehicle trajectories are obtained from the dataset [^3] Vehicle trajectory dataset: https://www.microsoft.com/en-us/research/publication/t-drive-trajectory-data-sample/, and the traffic flow data are sourced from the dataset444 California traffic flow dataset: https://pems.dot.ca.gov/. The main parameters used in the experiments are summarized in Table I [41, 8, 42, 43].
@@
The number of vehicles $(C)$ $5$
The number of UAVs $(U)$ $10$
@@
[…]
To demonstrate the effectiveness of the proposed DM-MAPPO algorithm, we compare it with several baseline algorithms, including Multi-Agent Soft Actor-Critic (MASAC), Greedy, and Random algorithms. The Greedy algorithm selects the edge server geographically closest to the vehicle for VT migration. To enhance exploration and avoid infeasible actions, a $\mathrm{10\%}$ probability is introduced for randomly selecting a server. The Random algorithm performs VT migration by randomly selecting the next edge server without any optimization or learned strategy.
VII-B Performance of LSTM-based Transformer Models on Workload Prediction
In Fig. 5, we show the predicted RSU workload curve generated by five different prediction algorithms. Among them, the proposed LSTM-based Transformer model yields a prediction curve that most closely aligns with the actual RSU workload. Figure 5 also shows the error rate of each prediction model in the high-frequency part of the traffic flow data, where the LSTM-based Transformer prediction error is always the smallest, outperforming the other models. As shown in Table II, the proposed model achieves the lowest Root Mean Square Error (RMSE) and Mean Absolute Error (MAE), indicating the smallest deviation between the predicted and actual values. Specifically, the proposed LSTM-based Transformer model achieves a prediction error rate of $\mathrm{6.814\%}$, highlighting its superior prediction performance. Moreover, the determination $R^{2}$ of the model approaches $1$, demonstrating strong explanatory power and the ability to capture most of the variance in the workload data. The LSTM-based Transformer model outperforms all baselines across all four evaluation metrics, This is because LSTM possesses the powerful capability of processing time-series data. By first leveraging LSTM to capture temporal dependencies and reduce high-frequency noise, followed by the self-attention mechanism of Transformer for fine-grained prediction refinement, the model delivers highly accurate results.
@@ VII-C4 Ablation study on the effectiveness of the dynamic mask module
VII-C5 Visual Analysis of multi-agent cooperative behavior
@@
To demonstrate the cooperation among agents in DM-MAPPO during the VT migration process, we develop a visualization platform to display the migration of each VT intuitively. As shown in Fig. 10, in time slot $t$, both vehicle $1$ and vehicle $3$ are traveling from north to south within the same region. As they approach the coverage boundary of RSU $7$, they need to migrate their VTs to edge servers with better channel quality, thus ensuring the immersion of VT services. For vehicle $1$, RSU $8$, and the satellite are identified as feasible edge server targets for VT migration. Given the higher communication cost associated with satellite links, vehicle $1$ opts to migrate its VTs to RSU $8$. Since the agents in DM-MAPPO share information such as workload information, vehicle $3$ migrating its VTs to RSU $8$ will overload RSU $8$, which may cause packet loss and affect the VT services of other vehicles. Therefore, for vehicle $3$, migrating its VTs to the satellite can maximize the overall benefit for all vehicles, which shows that our proposed DM-MAPPO can solve cooperative tasks well.
VII-C6 Visual analysis of the VT migration process
@@ VIII Conclusion
In this paper, we have addressed the challenges of optimizing multi-objective VT migration routing in 6G-enabled IoV networks. Specifically, we have proposed a workload prediction-based VT migration framework that leverages a heterogeneous edge server architecture comprising RSUs, UAVs, and the satellite. This architecture is particularly effective in mitigating the problem of low resource utilization in sparsely distributed regions. To enable accurate long-term workload prediction, we have proposed the LSTM-based Transformer model that integrates the LSTM module into the Transformer model. Furthermore, we have proposed the DM-MAPPO algorithm based on the dynamic mask module to facilitate efficient VT migration route planning in complex environments. The DM-MAPPO algorithm can greatly reduce the probability of infeasible actions and comprehensively optimize key performance metrics such as latency, packet loss rate, and regional workload balance. Simulation results have demonstrated that the LSTM-based Transformer model can reduce the error rate to $\mathrm{6.814\%}$, achieving accurate traffic flow prediction, and the performance of the DM-MAPPO algorithm has been improved by $\mathrm{73.41\%}$ compared with MAPPO. For future work, since our current framework considers all edge servers to be fully trustworthy, which may lead to the potential exposure of VT migration data and vehicle trajectories, we will explore the integration of privacy-preserving mechanisms into our framework.

### v2

@@ VI-B Algorithm Design
Algorithm 2 illustrates the pseudocode of the DM-MAPPO algorithm, whose computational complexity is $\mathcal{O}\left(End\times V\times t_{v}^{end}\times(M+end)\right)$, where $End$ represents the number of complete interactions of all agents with the environment, and endend represents the training epochs of each actor network.
@@ VII-A Parameter Settings
We first present the parameter settings for the experiments in the VT migration routing scenario. The scenario is sampled every $30\>\rm{s}$. Since all UAVs are designed to reach the next RSU for replenishment with the least energy consumption and in the shortest time, the trajectories of all UAVs are fixed, and the UAVs are set to patrol among RSUs at a speed of $3.7\>\rm{m/s}$[47]. For the network model, the mean of additive Gaussian white noise $\sigma$ is set to 0, with a variance of $10$ [5]. The vehicle trajectories are obtained from the dataset [^3] Vehicle trajectory dataset: https://www.microsoft.com/en-us/research/publication/t-drive-trajectory-data-sample/, and the traffic flow data are sourced from the dataset [^4] California traffic flow dataset: https://pems.dot.ca.gov/. The main parameters used in the experiments are summarized in Table I [10, 5, 25, 35].
@@
The number of vehicles $(C)$ 55
The number of UAVs $(U)$ 1010
@@
The weight for the latency objective ($w_{T}$) 22
The weight for the workload variance objective ($w_{s}$) 11
The weight for the packet loss rate objective ($w_{d}$) 11
To demonstrate the effectiveness of the proposed LSTM-based Transformer model, we compare it against four baseline models: LSTM, GRU, CNN-GRU, and CNN-LSTM. In both CNN-GRU and CNN-LSTM, the input data is initially processed using a one-dimensional convolution with a kernel size of $\mathrm{1}$, followed by prediction using GRU or LSTM, respectively [44, 31]. For the LSTM-based Transformer model, considering the factors of training time and prediction accuracy, we set the window of historical data input as $T=6$ to predict the traffic flow data in the next time slot $t+1$. In particular, for the LSTM prediction layer, the hidden unit in the LSTM module is set to $d_{h}=32$. For the Transformer prediction layer, the number of attention heads is set to $H=4$, and the dimension of query, key, and value vectors is set to $d_{k}=64$.
To demonstrate the effectiveness of the proposed DM-MAPPO algorithm, we compare it with several baseline algorithms, including Multi-Agent Soft Actor-Critic (MASAC), Greedy, and Random algorithms[41]. The Greedy algorithm selects the edge server geographically closest to the vehicle for VT migration. To enhance exploration and avoid infeasible actions, a $\mathrm{10\%}$ probability is introduced for randomly selecting a server. The Random algorithm performs VT migration by randomly selecting the next edge server without any optimization or learned strategy. To maintain environmental stability during the training of the DRL algorithm, the LSTM-based Transformer model should be trained before the DRL algorithm. This ensures that the workload of each edge server observed by the DRL algorithm remains consistent.
All experiments presented in this paper are conducted on a hardware platform equipped with an Intel Core i5 CPU and an MX450 GPU.
VII-B Performance of LSTM-based Transformer Model on Workload Prediction
In Fig. 5, we show the predicted RSU workload curve generated by five different prediction algorithms. Among them, the proposed LSTM-based Transformer model yields a prediction curve that most closely aligns with the actual RSU workload. Figure 5 also shows the error rate of each prediction model in the high-frequency part of the traffic flow data, where the LSTM-based Transformer prediction error is always the smallest, outperforming the other models. As shown in Table II, the proposed model achieves the lowest Root Mean Square Error (RMSE) and Mean Absolute Error (MAE), indicating the smallest deviation between the predicted and actual values. Specifically, the proposed LSTM-based Transformer model achieves a prediction error rate of $\mathrm{6.814\%}$, highlighting its superior prediction performance. Moreover, the determination $R^{2}$ of the model approaches $1$, demonstrating strong explanatory power and the ability to capture most of the variance in the workload data. The LSTM-based Transformer model outperforms all baselines across all four evaluation metrics. This is because the LSTM module possesses the powerful capability of processing time-series data. By first leveraging LSTM to capture temporal dependencies and reduce high-frequency noise, followed by the self-attention mechanism of Transformer for fine-grained prediction refinement, the model delivers highly accurate results, and its efficient training time of only $4$ minutes on the hardware platform demonstrates its strong potential for real-world applications.
@@ VII-C4 Ablation study on the effectiveness of the dynamic mask module
Figure 10: Algorithm performance comparison in scenarios with varying numbers of UAVs.
@@
VII-C5 Comparison of algorithm robustness under different numbers of edge servers
In practical scenarios, the total number of edge servers in an area is not fixed. This is mainly because UAVs may exit the region before vehicles reach their destinations, leading to dynamic variability in the number of available UAVs. Fortunately, with the introduction of a masking mechanism, the output of the action network can be expanded based on the currently known number of edge servers, with unused output entries being set to zero. The mask remains active until new UAVs enter the area. This approach provides a flexible interface for connecting intermittently arriving or departing UAVs to the DRL network, ensuring consistent integration under dynamic conditions.
The experimental setup is configured as follows: $5$ vehicles depart at time slot 0, while the remaining $2$ vehicles depart at different time slots, simulating dynamically entering vehicles. Among the edge servers, $5$ UAVs appear from the edge of the scenario at various times and begin providing VT services. In addition, $6$ UAVs move toward different edges of the area at different times and exit the region. As shown in Fig. 10, DM-MAPPO still maintains the best performance under these experimental conditions, achieving the lowest latency and packet loss rate. This demonstrates that DM-MAPPO possesses strong robustness, enabling it to effectively adapt to highly dynamic environments, and thus making it well-suited for real-world applications.
VII-C6 Visual analysis of multi-agent cooperative behavior
To demonstrate the cooperation among agents in DM-MAPPO during the VT migration process, we develop a visualization platform to display the migration of each VT intuitively. Typically, the number of vehicles exceeds that of RSUs. However, the edge servers providing services to these five vehicles are also capable of serving other vehicles. For a simple and clear presentation, we have exclusively displayed the behavioral trajectories of these $5$ vehicles controlled by the DM-MAPPO algorithm. As shown in Fig.11, in time slot $t$, both vehicle $1$ and vehicle $3$ are traveling from north to south within the same region. As they approach the coverage boundary of RSU $7$, they need to migrate their VTs to edge servers with better channel quality, thus ensuring the immersion of VT services. For the vehicle $1$, RSU $8$, and the satellite are identified as feasible edge server targets for VT migration. Given the higher communication cost associated with satellite links, vehicle $1$ opts to migrate its VTs to RSU $8$. Since the agents in DM-MAPPO share information such as workload information, vehicle $3$ migrating its VTs to RSU $8$ will overload RSU $8$, which may cause packet loss and affect the VT services of other vehicles. Therefore, for vehicle $3$, migrating its VTs to the satellite can maximize the overall benefit for all vehicles, which shows that our proposed DM-MAPPO can solve cooperative tasks well.
VII-C7 Visual analysis of the VT migration process
@@ VIII Conclusion
In this paper, we have addressed the challenges of optimizing multi-objective VT migration routing in 6G-enabled IoV networks. Specifically, we have proposed a workload prediction-based VT migration framework that leverages a heterogeneous edge server architecture comprising RSUs, UAVs, and the satellite. This architecture is particularly effective in mitigating the problem of low resource utilization in sparsely distributed regions. To enable accurate long-term workload prediction, we have proposed the LSTM-based Transformer model that integrates the LSTM module into the Transformer model. Furthermore, we have proposed the DM-MAPPO algorithm based on the dynamic mask module to facilitate efficient VT migration route planning in complex environments. The DM-MAPPO algorithm can greatly reduce the probability of infeasible actions and comprehensively optimize key performance metrics such as latency, packet loss rate, and regional workload balance. Simulation results have demonstrated that the LSTM-based Transformer model can reduce the error rate to $\mathrm{6.814\%}$, achieving accurate traffic flow prediction, and the performance of the DM-MAPPO algorithm has been improved by $\mathrm{73.41\%}$ compared with MAPPO. The proposed framework operates under the idealistic assumptions of fully trustworthy edge servers and simplified channel models. However, these assumptions may lead to risks such as the exposure of VT migration data and vehicle trajectories, potentially hindering the framework’s practical application. Therefore, for future work, we will explore the channel model with line-of-sight and non-line-of-sight natures and incorporate privacy-preserving mechanisms into our framework[48, 3].